A tour operating firm plans to take tourists between Addis Ababa and Jimma. its estimated cost function given by C =100 +50 N+4N2 (Where N denotes the number of passengers per day)


       A. State the average cost function.

       B. State the marginal cost of the average cost function.

       C. Find the number of passengers per day that minimize average cost.

       D. What is the minimum average cost at the optimal level of passenger?

       E. What will be the total cost at the optimal level?




1
Expert's answer
2022-01-16T17:47:38-0500

(a) Average cost function ="\frac{100+50n+4n^2}{n}" =

"\frac{100}{n}+50+4n"

(b) C(n) ="100+50n+4n^{2}"

"C^{'} (n) =50+8n"

(c) marginal cost function =average cost function

"8n+50=\frac{100}{n}+50+4n"

"4n^2=100"

n=5 passengers per day

(d) The minimum average cost "=\frac{100}{5}+50+4(5)"

"\therefore" minimum average cost="90\$"

(e) The total cost at optimal level ="100+50(5)+4(5^{2})" "=450\$"